OpenStudy (anonymous):

What is the solution to the rational equation x over x squared minus 9 plus 1 over x minus 3 equals 1 over 4x minus 12 ? I honestly dont know how to do this but i'll give a medal to anyone who can work it out

OpenStudy (blurbendy):

x / (x^2 - 9) + 1 / (x - 3) = 1 / (4x - 12) x = -9/7 Do you need help arriving to that solution?

OpenStudy (blurbendy):

Remember, you can message me for help if you don't get an adequate response quickly enough.